OpenStudy (anonymous):

What is the value of log81 3?

OpenStudy (anonymous):

\[\log _{81}3\]

OpenStudy (mathmale):

There are various ways in which to do this problem. One of the easier approaches is to re-write \[y=\log_{81} 3\] as y=\[y=\log_{3^4} 3\]. All I did here was to replace the base 81 with its equivalent as a power of 3.

OpenStudy (jdoe0001):

\(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ \quad \\ log_{\color{red}{ 81}}{\color{blue}{ 3}}=\square \implies ?\)

OpenStudy (anonymous):

81^y=3

OpenStudy (jdoe0001):

hmm yes

OpenStudy (jdoe0001):

\(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ \quad \\ log_{\color{red}{ 81}}{\color{blue}{ 3}}=\square \implies {\color{red}{ 81}}^y=3 \\ \quad \\ {\color{brown}{ 81\to 3\cdot 3\cdot 3\cdot 3\to 3^2\cdot 3^2\to (3^2)^2\to 3^4}}\qquad thus \\ \quad \\ 81^y=3\implies (3^4)^y=3^1\implies 3^{{\color{blue}{ 4y}}}=3^{\color{blue}{ 1}}\)

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \log_{81}3 }\) --> \(\Large\color{blue}{ \log_{81}81^{-1/4} }\)

OpenStudy (anonymous):

So 3^4y = 3^1 is my answer?

OpenStudy (solomonzelman):

( damn, Greece missed again.... )

OpenStudy (jdoe0001):

hheh

OpenStudy (jdoe0001):

\(\bf (3^4)^y=3^1\implies 3^{{\color{blue}{ 4y}}}=3^{\color{blue}{ 1}}\implies {\color{blue}{ 4y}}={\color{blue}{ 1}}\)

OpenStudy (mathmale):

Let Solomon finish, please. He / you converted the '3' on the extreme right in to 81^(1/4), which was a clever approach. But what about the 'y' on the left side?

OpenStudy (solomonzelman):

I can't finish, I am watching a match of Greece and Puerto Rice. I am like here and there....

OpenStudy (solomonzelman):

Costo rico

OpenStudy (dangerousjesse):

Mm rice

OpenStudy (solomonzelman):

OMG.... Greece couldn't finish the attack!!

OpenStudy (solomonzelman):

I am out, I am commenting about soccer and shouldn't be doing that.

OpenStudy (anonymous):

so my answer 4y=1

OpenStudy (johnweldon1993):

\[\large log_{81} 3 = x \Rightarrow 81^x = 3\] Take the log of both sides We know that \(\huge \ln(a)^b = b \times \ln(a)\) so... \[\large x \times \ln (81) = \ln(3)\] Divide both sides by \(\large \ln(81)\) \[\large x = \frac{\ln(3)}{\ln(81)}\] We find that \[\large x = 4\]

OpenStudy (mathmale):

Note that both Solomon and Joe are attempting to convert the given problem so that both the base and the quantity of which the log is taken have the same base. If Solomon is unavailable, then Joe: would you like to guide Skittle towards finding her own answer?